OpenStudy (anonymous):

how do i factor x^4+5x^2+6=?

OpenStudy (cwrw238):

let x^2 = t = t^2 + 5t + 6 now you need 2 numbers a and b where ab = +6 and a + b = +5 the factors will be of the form (t + a)(t + b)

OpenStudy (cwrw238):

can you figure it out?

OpenStudy (cwrw238):

the 2 numbers are + 2 and +3 2*3 = 6 and 2 + 3 = 5

OpenStudy (anonymous):

yes thank you very much, you are a life savor!

OpenStudy (cwrw238):

yw - finally substitute back x^2 for t

OpenStudy (anonymous):

is that ur cat in pic?? he/she is angry !!!

OpenStudy (cwrw238):

lol!! - yes she's mine - she does look that way . Normally she's very friendly - except when there's other cats around.